OpenStudy (anonymous):

2x+3y=26, 2y+z=19 and x+2z=29 what is the value of x+y+y?

OpenStudy (anonymous):

So whats the question exactly ???

OpenStudy (anonymous):

2x+3y=26, 2y+z=19 and x+2z=29 what is the value of x+y+y?

OpenStudy (anonymous):

substitute all the values.. and then add the constants ... there u r wid ur answer ... !

OpenStudy (zzr0ck3r):

do you know linear algebra?

OpenStudy (zzr0ck3r):

you can out this in a matrix and row reduce to get x=7, y=4, z=11

OpenStudy (zzr0ck3r):

so 15

OpenStudy (anonymous):

no bro its 22 ans

OpenStudy (zzr0ck3r):

then you meant x+y+z but you wrote x+y+y

OpenStudy (zzr0ck3r):

x+y+z is indeed 22, but x+y+y is not